# Accelerating Quantum Monte Carlo Calculations with Set-Equivariant Architectures and Transfer Learning

### v2

@@ Abstract
Machine-learning (ML) ansätze have greatly expanded the accuracy and reach of variational quantum Monte Carlo (QMC) calculations, in particular when exploring the manifold quantum phenomena exhibited by spin systems. However, the scalability of QMC is still compromised by several other bottlenecks, and specifically those related to the actual evaluation of observables based on random deviates that lies at the core of the approach. Here we show how the set-transformer architecture can be used to dramatically accelerate or even bypass that step, especially for time-consuming operators such as powers of the magnetization. We illustrate the procedure with a range of examples of increasing complexity, from the classical Ising model to quantum systems with long-range interactions, and comprising both regressions (to predict observables) and classifications (to detect phase transitions). Moreover, we show how transfer learning can be leveraged to reduce the training cost by reusing knowledge from different systems and smaller system sizes.
@@ 1 Introduction
The explosion of machine learning (ML) in the last two decades has had an extensive and deep impact on condensed matter physics and materials science, with every new advance in architectures and training quickly finding applications in those disciplines [1, 2]. A good illustration is provided by the introduction of advanced regression models as interatomic potentials with ab-initio quality, starting with basic multilayer perceptrons [3] and moving towards sophisticated equivariant graph neural networks [4]. Classification models have been equally prolific, with abundant examples of applications for the analysis of spectroscopic data, as a close analogue of image classification tasks [5]. Likewise, generative models have found many natural uses, some closely related to more mainstream tasks —like the generation of candidate crystalline structures [6], analogous to image or video generation— and some field-specific —such the creation of Boltzmann generators based on normalizing flows to efficiently sample the phase space of classical systems [7].
In the last years, ML strategies have also been deployed to assist in the QMC-based exploration of the rich physics of systems beyond the reach of analytical solutions. For instance, classifiers have been employed to detect phase transitions in problematic many-fermion systems [8, 9] and in a more general setting [10]. Moreover, very general functional forms for regression and methods to train them efficiently have appeared on the scene [11], which can be used as flexible variational ansätze for a search of the ground state through parametric minimization. The process is similar to —and can therefore reuse part of the machinery developed for— the stochastic minimization of a loss function when training a typical ML regression model, albeit with some important modifications. Firstly, the geometry of the quantum ground state problem favors stochastic reconfiguration [12, 13], a minimization method harnessing more information about the target function than standard ML minimizers like ADAM [14]. More importantly, the quantity being minimized is not a metric of the difference with the ground truth for a predefined set of inputs but a physical quantity. Hence, the concerns about data availability that might be present in an ML regression setting are replaced by the complexity of sampling from the ground-state probability density function and evaluating the expected value of the Hamiltonian and other operators based on those samples. That complexity remains one of the causes of bottlenecks in variational QMC.
@@
In this paper we present an accelerated approach to the calculation of expectation values by making more efficient use of the samples drawn from $\left|\psi\right|^{2}$. A critical feature of such approaches is that the estimate of an expectation value cannot depend on the order in which the $\left\lvert\nu\right\rangle$ are drawn from $\left|\psi\right|^{2}$ or fed to the model, i.e., it must be permutation invariant. Moreover, they must be able to work on collections of $\left\lvert\nu\right\rangle$ with different cardinality. Those are the characteristics of a set-invariant architecture. We have therefore based our strategy on an architecture developed as a universal, efficient approximation to set-equivariant functions, the set transformer [17]. We use it as the backbone of permutation-invariant regression models. We also show how to build a classifier based on the set transformer to detect phase transitions. The calculation of each expectation value using the training architecture is three to four orders of magnitude faster than the direct approach. In addition to analyzing the performance of these models, we discuss how to build them more efficiently by amortizing the training cost over ranges of Hamiltonian parameters, numbers of samples and system sizes: for instance, after pretraining for a chain of $50$ spins, extending the model to $100$ and $150$ spins only requires an additional $70\%$ computational effort.
@@ 2.1 The Set Transformer Architecture
The deep sets theorem [18] states that a universal form of a permutation-invariant $f_{\mathcal{O}}\left(\mathcal{S}\right)$ is
$f_{\mathcal{O}}\left(\mathcal{S}\right)=G\left[\sum\limits_{\nu\in\mathcal{S}}g\left(\lvert\nu\rangle\right)\right].$
@@
Hence, in principle a permutation-invariant regression model could be built by choosing and training sufficiently general ML models to take the roles of $g$ and $G$. This has been successfully exploited in ML force fields [19], where the inputs are atomic coordinates and types, the output is the system’s potential energy, and $G$ is taken as the identity to create an additive architecture. However, in a more general setting the scalability and expressive power of actual realizations of this construction based on finite models are quite limited [20], and the same applies to slight variations using set-pooling functions other than the sum.
@@
A more practically oriented approach to permutation invariance that still results in an universal approximator of set-invariant functions is the set transformer [17], derived from the extremely successful transformer architecture [21]. The set transformer can be described starting from a conventional multihead dot-product attention block with LayerNorm normalization [22] but without any positional encoding [23] or dropout [24]. We will denote that block as $\operatorname{MAB}\left(\bm{X},\bm{Y}\right)$, where $\bm{X}$ and $\bm{Y}$ are two arrays of $n$ $d$-dimensional tokens each, with $\bm{X}$ being used to generate the queries and $\bm{Y}$ being used to generate the keys and values of the attention mechanism. This basic $\operatorname{MAB}$ block is then used to build two different kinds of higher-level components: set-attention blocks ($\operatorname{SAB}$) and pooling-by-multihead-attention ($\operatorname{PMA}_{k}$) blocks. While $\operatorname{PMA}_{k}$ blocks are set invariant, $\operatorname{SAB}$s are set equivariant, i.e., $\operatorname{SAB}\left[\tau\left(\bm{X}\right)\right]=\tau\left[\operatorname{SAB}\left[\bm{X}\right]\right]\;\forall\tau\in\operatorname{Sym}\left(\bm{X}\right)$. A full set transformer contains an encoder pipeline in the form of a stack of set-equivariant blocks followed by a decoder stage producing the final output. In our models we design the decoder stage to be set invariant, thus making the whole transformer invariant as well.
@@
A set-equivariant block can be straightforwardly built by turning the multihead attention block into a self-attention block by short-circuiting its inputs $\operatorname{SAB}\left(\bm{X}\right)=\operatorname{MAB}\left(\bm{X},\bm{X}\right)$; in fact, it is this well-known property of the dot-product self attention that motivates the introduction of positional encodings in applications where neighborhood relations between input tokens are important and therefore set equivariance is undesirable. However, the scalability of this simple $\operatorname{SAB}$ is limited and a more performance-oriented alternative is the induced set attention block $\operatorname{ISAB}_{m}\left(\bm{X}\right)=\operatorname{MAB}\left[X,\operatorname{MAB}\left(\bm{I},\bm{X}\right)\right]$, where $\bm{I}$ is a set of $m<n$ trainable inducing points.
@@
For the system sizes employed in this work, we do not need to resort to the ISAB. Our set transformers are built entirely our of self-attention and pooling-by-self-attention blocks. The whole set of building blocks is schematically represented in Figure 1; following the original recipe [17], in this work we employ implementations of the set transformer where each processing component is simply a non-linear unit, more specifically a ReLU [25].
@@
Figure 1: Structure of the main building blocks of the Set Transformer. The Multihead Attention Block (MAB) is the core component, defined as a standard multihead dot-product attention with LayerNorm and ReLU, but without positional encoding or dropout. The Induced Set Attention Block (ISAB) is an alternative to self-attention that preserves the set equivariance but improves scaling; it is constructed as $\operatorname{ISAB}_{m}\left(\bm{X}\right)=\operatorname{MAB}\left[\bm{X},\operatorname{MAB}\left(\bm{I},\bm{X}\right)\right]$ with $m<n$ trainable inducing points $\bm{I}$. The Pooling by Multihead Attention (PMA) block defines the set-invariant decoder stage, specified as $\operatorname{PMA}_{k}\left(\bm{X}\right)=\operatorname{MAB}\left[\bm{S},\operatorname{FF}\left(\bm{X}\right)\right]$, where $\bm{S}$ are $k$ trainable seed vectors and $FF$ is a feed-forward block.
@@
A set-transformer-based approach is model-agnostic, and can be applied in other kinds of QMC. The only requirement is access to configurations drawn from a sampling measure, be they fermionic samples, occupation snapshots, Hubbard–Stratonovich slices, or equal-time Green matrices. For instance, in Ref. [8], Broecker et al. demonstrated that a CNN trained on equal-time Green matrices produced by auxiliary-field QMC can locate fermionic phase transitions even when conventional sign-reweighted estimators fail. Their key idea is to “state function” directly from QMC configurations, thereby bypassing explicit sign reweighting in the final prediction. This complements our set-equivariant approach: a set transformer can be trained with exactly the same sampled objects and perform phase classification without computing observables, thus inheriting the same practical insensitivity to the sign problem for classification tasks.
@@ 2.2 Our Regression and Classification Models
We use two specific set transformer models, one for regression and one for classification. Our classifier contains an encoder stage made up of a $5$-head $\operatorname{SAB}$ followed by a $10$-head SAB, a decoder stage comprising a $5$-head, single-seed $\operatorname{PMA}$ followed by two $10$-head $\operatorname{SAB}$s with a LayerNorm between them, a linear layer to generate the logits, and a SoftMax layer to generate the final output. We do not include any $\operatorname{ISAB}$ block since the quadratic scaling of attention does not become a problem for these system sizes. When we apply this model to label phases as antiferromagnetic, ferromagnetic or paramagnetic, we obtain better results by using a sequence of two binary classifiers than by resorting to a direct ternary classifier. The first of those classifiers sets aside paramagnetic states; the remaining ones are passed on to the second classifier, which then discriminates between ferromagnetic and antiferromagnetic states.
@@ 2.3 Hamiltonian, Ansatz and Ground-State Search
where $\left|i-j\right|$ must be understood in the context of the periodic boundary conditions and $\eta=1+\sum\limits_{j=1}^{N}j^{-\alpha}$. Thus, the three key parameters of the model are the size, i.e. the number os spins, $N$; the ratio $J/h$ that controls the strength of the interactions between spins, and $\alpha$, which determines how fast those interactions decay with distance. The role of the Kac renormalization factor, $\eta$, is to render the model extensive with respect to $N$ [26, 27] .
We choose this model because it provides a tunable interpolation between two paradigmatic and exactly solvable limits: the transverse-field Ising model with nearest-neighbor interactions, recovered in the limit $\alpha\to\infty$, and the fully-connected Lipkin-Meshkov-Glick (LMG) model at $\alpha=0$. The former is not only analytically tractable but also serves as the standard benchmark for quantum simulations using both machine learning and other numerical techniques. The latter is also exactly solvable and well understood, offering a complementary point of comparison. By continuously varying $\alpha$, the model exhibits a rich phenomenology arising from the interplay between interaction range and quantum fluctuations. For $\alpha>\alpha^{*}$, the system behaves effectively as short-range, while for $\alpha\in(0,1)$ strong long-range effects emerge, with $\eta$ diverging as a function of system size $N$, thus becoming essential to recover the thermodynamic limit as $N\to\infty$. In between those, a weak long-range regime appears. Moreover, the model has recently been studied across the full range of $\alpha$ using a variety of methods including QNNs [28, 29, 30, 31], enabling us to benchmark our results and quantify the numerical speedup achieved by our approach.
To find the ground state of the Hamiltonian in (2) we optimize the parameters of a neural-quantum-state ansatz. In particular, we choose a flexible functional form based on the vision transformer (ViT [32, 11]), another widely used specialization of the transformer architecture. In our model, the projection of state $\left\lvert\psi\right\rangle$, represented by the ansatz, on element $\lvert s\rangle$ of the computational basis is determined through the following steps:
@@
    The processed tokens are normalized by a LayerNorm [22] and fed into an $\operatorname{MAB}$ with $H$ heads. The attention models the interaction between tokens. To make this interaction respect translational symmetry, the attention is parameterized by a circulant matrix.
@@
As in the case of the set transformer, within each core block, two ResNet-style shortcuts [33] allow the network to be trained to bypass the mixing and processing steps if necessary to improve performance and therefore avoid the drawbacks of increased network depth. To enforce translation symmetry also at the intra-token level, the sequence above is repeated $d$ times with the input shifted by one position each time with respect to the previous one, so that every spin occupies all possible positions within each token, and the final output is taken as the average of those $d$ runs.

### v3

@@ Abstract
Machine-learning (ML) ansätze have greatly expanded the accuracy and reach of variational quantum Monte Carlo (QMC) calculations, in particular when exploring the manifold quantum phenomena exhibited by spin systems. However, the scalability of QMC is still compromised by several other bottlenecks, and specifically those related to the actual evaluation of observables based on random deviates that lies at the core of the approach. Here we show how the set-transformer architecture can be used to dramatically accelerate or even bypass that step, especially for time-consuming operators such as powers of the magnetization. We illustrate the procedure with a range of examples structured around quantum spin systems with long-range interactions, and comprising both regressions (to predict observables) and classifications (to detect phase transitions). Moreover, we show how transfer learning can be leveraged to reduce the training cost by reusing knowledge from different systems and smaller system sizes.
@@ 1 Introduction
The explosion of machine learning (ML) in the last two decades has had an extensive and deep impact on condensed matter physics and materials science, with every new advance in architectures and training quickly finding applications in those disciplines [Carleo2019, Malica2025]. A good illustration is provided by the introduction of advanced regression models as interatomic potentials with ab-initio quality, starting with basic multilayer perceptrons [behler] and moving towards sophisticated equivariant graph neural networks [nequip]. Classification models have been equally prolific, with abundant examples of applications for the analysis of spectroscopic data, as a close analogue of image classification tasks [classifier_examples]. Likewise, generative models have found many natural uses, some closely related to more mainstream tasks —like the generation of candidate crystalline structures [generative], analogous to image or video generation— and some field-specific —such the creation of Boltzmann generators based on normalizing flows to efficiently sample the phase space of classical systems [normalizing_flows].
In the last years, ML strategies have also been deployed to assist in the QMC-based exploration of the rich physics of systems beyond the reach of analytical solutions. For instance, classifiers have been employed to detect phase transitions in problematic many-fermion systems [Broecker2017, Chng2017] and in a more general setting [ghosh2025learningphasesquantummonte]. Moreover, very general functional forms for regression and methods to train them efficiently have appeared on the scene [Viteritti2023], which can be used as flexible variational ansätze for a ground-state search through parametric minimization. The process is similar to —and can therefore reuse part of the machinery developed for— the stochastic minimization of a loss function when training a typical ML regression model, albeit with some important modifications. Firstly, the geometry of the quantum ground state problem favors stochastic reconfiguration [PhysRevLett.80.4558, PhysRevB.64.024512], a minimization method harnessing more information about the target function than standard ML minimizers like ADAM [ADAM]. More importantly, the quantity being minimized is not a metric of the difference with the ground truth for a predefined set of inputs, but a physical quantity. Hence, the concerns about data availability that might be present in an ML regression setting are replaced by the complexity of sampling from the ground-state probability density function and evaluating the expected value of the Hamiltonian and other operators based on those samples. That complexity remains one of the causes of bottlenecks in variational QMC.
@@
In this paper we present an accelerated approach to the calculation of expectation values by making more efficient use of the samples drawn from $\left|\psi\right|^{2}$. A critical feature of such approaches is that the estimate of an expectation value cannot depend on the order in which the $\left\lvert\nu\right\rangle$ are drawn from $\left|\psi\right|^{2}$ or fed to the model, i.e., it must be permutation invariant. Moreover, they must be able to work on collections of $\left\lvert\nu\right\rangle$ with different cardinality. Those are the characteristics of a set-invariant architecture. We have therefore based our strategy on an architecture developed as a universal, efficient approximation to set-equivariant functions, the set transformer [set_transformer]. We use it as the backbone of permutation-invariant regression models. We also show how to build a classifier based on the set transformer to detect phase transitions. The calculation of each expectation value using the training architecture is three to four orders of magnitude faster than the direct approach. In addition to analyzing the performance of these models, we discuss how to build them more efficiently by amortizing the training cost over ranges of Hamiltonian parameters, numbers of samples, and system sizes: for instance, after pretraining for a chain of $50$ spins, extending the model to $100$ and $150$ spins only requires an additional $70\%$ computational effort.
@@ 2.1 The Set-Transformer Architecture
The deep-sets theorem [deep_sets] states that a universal form of a permutation-invariant $f_{\mathcal{O}}\left(\mathcal{S}\right)$ is
$f_{\mathcal{O}}\left(\mathcal{S}\right)=G\left[\sum\limits_{\nu\in\mathcal{S}}g\left(\lvert\nu\rangle\right)\right].$ (2)
@@
In principle, a permutation-invariant regression model could hence be built by choosing and training sufficiently general ML models to take the roles of $g$ and $G$. When dealing with sampled spin configurations $\left\lvert\nu\right\rangle\in\mathcal{C}$, these deep-sets (or set-pooling) architectures take the form of conventional (e.g. fully connected or convolutional) models that are applied to each $\left\lvert\nu\right\rangle$ independently, and whose results are summed or averaged to produce the final output. This scheme has also been successfully exploited in ML force fields [mlff_uncertainties], where the inputs are atomic coordinates and types, the output is the system’s potential energy, and $G$ is taken as the identity to create an additive architecture. However, in a more general setting the scalability and expressive power of actual realizations of this construction based on finite models have been shown to be limited [deep_sets_limitations], and the same applies to slight variations using set-pooling functions other than the sum.
@@
A more practically oriented approach to permutation invariance that still results in an universal approximator of set-invariant functions is the set transformer [set_transformer], derived from the extremely successful transformer architecture [attention]. The set transformer can be described starting from a conventional multihead dot-product attention block with LayerNorm normalization [LayerNorm] but without any positional encoding [Shaw2018] or dropout [dropout]. We will denote that block as $\operatorname{MAB}\left(\bm{X},\bm{Y}\right)$, where $\bm{X}$ and $\bm{Y}$ are two arrays of $n$ $d$-dimensional tokens each, with $\bm{X}$ being used to generate the queries and $\bm{Y}$ being used to generate the keys and values of the attention mechanism. This basic $\operatorname{MAB}$ block is then used to build two different kinds of higher-level components: set-attention blocks ($\operatorname{SAB}$) and pooling-by-multihead-attention ($\operatorname{PMA}_{k}$) blocks. While $\operatorname{PMA}_{k}$ blocks are set invariant, $\operatorname{SAB}$s are set equivariant, i.e., $\operatorname{SAB}\left[\tau\left(\bm{X}\right)\right]=\tau\left[\operatorname{SAB}\left[\bm{X}\right]\right]\;\forall\tau\in\operatorname{Sym}\left(\bm{X}\right)$. A full set transformer contains an encoder pipeline in the form of a stack of set-equivariant blocks followed by a decoder stage producing the final output. In our models we design the decoder stage to be set invariant, thus making the whole transformer invariant as well.
@@
A set-equivariant block can be straightforwardly built by turning the multihead attention block into a self-attention block by short-circuiting its inputs $\operatorname{SAB}\left(\bm{X}\right)=\operatorname{MAB}\left(\bm{X},\bm{X}\right)$]; in fact, it is this well-known property of the dot-product self attention that motivates the introduction of positional encodings in applications where neighborhood relations between input tokens are important and therefore set equivariance is undesirable. However, the scalability of this simple $\operatorname{SAB}$ is limited and a more performance-oriented alternative is the induced set-attention block $\operatorname{ISAB}_{m}\left(\bm{X}\right)=\operatorname{MAB}\left[\bm{X},\operatorname{MAB}\left(\bm{I},\bm{X}\right)\right]$, where $\bm{I}$ is a set of $m<n$ trainable inducing points.
@@
For the system sizes employed in this work, we do not need to resort to the ISAB. Our set transformers are built entirely out of self-attention and pooling-by-self-attention blocks. The whole set of building blocks is schematically represented in Figure 1; following the original recipe [set_transformer], in this work we employ implementations of the set transformer where each processing component is simply a non-linear unit, more specifically a ReLU [ReLU].
@@
Figure 1: Structure of the main building blocks of the set transformer. The multihead attention block (MAB) is the core component, defined as a standard multihead dot-product attention with LayerNorm and ReLU, but without positional encoding or dropout. The induced set-attention Block (ISAB) is an alternative to self-attention that preserves the set equivariance but improves scaling; it is constructed as $\operatorname{ISAB}_{m}\left(\bm{X}\right)=\operatorname{MAB}\left[\bm{X},\operatorname{MAB}\left(\bm{I},\bm{X}\right)\right]$ with $m<n$ trainable inducing points $\bm{I}$. The pooling by multihead attention (PMA) block defines the set-invariant decoder stage, specified as $\operatorname{PMA}_{k}\left(\bm{X}\right)=\operatorname{MAB}\left[\bm{S},\operatorname{FF}\left(\bm{X}\right)\right]$, where $\bm{S}$ are $k$ trainable seed vectors and $FF$ is a feed-forward block.
@@
A set-transformer-based approach is model-agnostic, and can be applied in other kinds of QMC. The only requirement is access to configurations drawn from a sampling measure, be they fermionic samples, occupation snapshots, Hubbard–Stratonovich slices, or equal-time Green matrices. For instance, in Ref. [Broecker2017], Broecker et al. demonstrated that a CNN trained on equal-time Green matrices produced by auxiliary-field QMC can locate fermionic phase transitions even when conventional sign-reweighted estimators fail. Their key idea is to approximate the “state function” directly from QMC configurations, thereby bypassing explicit sign reweighting in the final prediction. This complements our set-equivariant approach: a set transformer can be trained with exactly the same sampled objects and perform phase classification without computing observables, thus inheriting the same practical insensitivity to the sign problem for classification tasks.
@@ 2.2 Our Regression and Classification Models
We use two specific set-transformer models, one for regression and one for classification. Our classifier contains an encoder stage made up of a $5$-head $\operatorname{SAB}$ followed by a $10$-head SAB, a decoder stage comprising a $5$-head, single-seed $\operatorname{PMA}$ followed by two $10$-head $\operatorname{SAB}$s with a LayerNorm between them, a linear layer to generate the logits, and a SoftMax layer to generate the final output. We do not include any $\operatorname{ISAB}$ block since the quadratic scaling of attention does not become a problem for these system sizes. When we apply this model to label phases as antiferromagnetic, ferromagnetic or paramagnetic, we obtain better results by using a sequence of two binary classifiers than by resorting to a direct ternary classifier. The first of those classifiers sets aside paramagnetic states; the remaining ones are passed on to the second classifier, which then discriminates between ferromagnetic and antiferromagnetic states.
@@ 2.3 Hamiltonian, Ansatz and Ground-State Search
where $\left|i-j\right|$ must be understood in the context of the periodic boundary conditions and $\eta=1+\sum\limits_{j=1}^{N}j^{-\alpha}$. Thus, the three key parameters of the model are the size, i.e. the number os spins, $N$; the ratio $J/h$ that controls the strength of the interactions between spins, and $\alpha$, which determines how fast those interactions decay with distance. The role of the Kac renormalization factor, $\eta$, is to render the model extensive with respect to $N$ [campa2009statistical, romanroche2023] .
We choose this model because it provides a tunable interpolation between two paradigmatic and exactly solvable limits: the transverse-field Ising model with nearest-neighbor interactions, recovered in the limit $\alpha\to\infty$, and the fully-connected Lipkin-Meshkov-Glick (LMG) model at $\alpha=0$. The former is not only analytically tractable but also serves as the standard benchmark for quantum simulations using both machine learning and other numerical techniques. The latter is also exactly solvable and well understood, offering a complementary point of comparison. By continuously varying $\alpha$, the model exhibits a rich phenomenology arising from the interplay between interaction range and quantum fluctuations. For $\alpha>\alpha^{*}$, the system behaves effectively as short-range, while for $\alpha\in(0,1)$ strong long-range effects emerge, with $\eta$ diverging as a function of system size $N$, thus becoming essential to recover the thermodynamic limit as $N\to\infty$. In between those, a weak long-range regime appears. Moreover, the model has recently been studied across the full range of $\alpha$ using a variety of methods including QNNs [koziol2021quantumcritical, Kim2024, PhysRevB.110.205147, HerraizLpez2025], enabling us to benchmark our results and quantify the numerical speedup achieved by our approach.
To find the ground state of the Hamiltonian in (3), we optimize the parameters of a neural-quantum-state ansatz. In particular, we choose a flexible functional form based on the vision transformer (ViT [vision_transformer, Viteritti2023]), another widely used specialization of the transformer architecture. In our model, the projection of state $\left\lvert\psi\right\rangle$, represented by the ansatz, on element $\lvert s\rangle$ of the computational basis is determined through the following steps:
@@
    The processed tokens are normalized by a LayerNorm [LayerNorm] and fed into an $\operatorname{MAB}$ with $H$ heads. The attention models the interaction between tokens. To make this interaction respect translational symmetry, the attention is parameterized by a circulant matrix.
@@
As in the case of the set transformer, within each core block, two ResNet-style shortcuts [ResNet] allow the network to be trained to bypass the mixing and processing steps if necessary to improve performance and therefore avoid the drawbacks of increased network depth. To enforce translation symmetry also at the intra-token level, the sequence above is repeated $d$ times with the input shifted by one position each time with respect to the previous one, so that every spin occupies all possible positions within each token, and the final output is taken as the average of those $d$ runs.